OpenStudy (anonymous):

Metal for reward! A TV station’s weather radar can detect precipitation in a circular region having a diameter of 100 mi. Find the area of the region. a. 31,400 mi2 c. 15,700 mi2 b. 7,850 mi2 d. 3,140 mi2

OpenStudy (austinl):

\[A=\pi\times r^2\]

OpenStudy (anonymous):

it doesn't come out right

OpenStudy (anonymous):

\[r = \frac{1}{2}d\]

OpenStudy (anonymous):

Wait, do you have a picture of the problem? None of these answers would be correct.

OpenStudy (anonymous):

no, i don't

OpenStudy (anonymous):

Oh I get it now. First find what the radius is \(\left(r = \frac{1}{2}d\right)\), then plug it into the area formula \(\left(A=\pi r^2\right)\). Then, they want you to approximate \(\pi\) as 3.14 and multiply out your area. Then choose the closest answer.

OpenStudy (anonymous):

yes, what would it be?

OpenStudy (anonymous):

Well, how far can you get on your own?

OpenStudy (anonymous):

iI got 200 because I don't know how to solve for the given price

OpenStudy (anonymous):

No, the \(diameter\) is 100. The radius is one-half the diameter. So, what's the radius?